1. How much longer in percentage is a one mile race than 1500 m race? [2]
2. A typical atom has a diameter of about 4.56 x10 -10 m
a) What is the diameter in Inches and µm [2]
b) Approximately how many atoms are there along a 1.0 cm line [2]
1
Expert's answer
2020-05-07T10:08:27-0400

1)


"1-\frac{1500}{1609}=0.0677=6.77\%"

2) a)


"d=4.56 \cdot10^{ -10}\cdot10^6=4.56\cdot10^{ -4}\mu m"

"d=4.56 \cdot10^{ -10}\cdot39.37=1.80\cdot10^{ -8}\ in"

b)

"n=\frac{0.01}{4.56 \cdot10^{ -10}}=2.19\cdot10^{ 7}"